# Semi-supervised Semantic Segmentation for Remote Sensing Images via Multi-scale Uncertainty Consistency and Cross-Teacher-Student Attention

### v2

@@ II-A RS semantic segmentation
Semantic segmentation is to segment an image into regional blocks with certain semantic meanings [22], and recognize the specific semantic category of each regional block respectively. It commonly labels the image pixel by pixel [23, 24, 25, 26, 27]. Early semantic segmentation models including FCN [28], SegNet [29], and U-Net [30], achieved remarkable performance in traditional semantic segmentation tasks. With the success of Transformers in vision in recent years, more and more semantic segmentation models are adopting the Transformers architecture[31]. Complex backgrounds as well as resolution variations are the main challenges in semantic segmentation of RS images. In RS images, large intra-class and small inter-class variations between objects make image feature characterization difficult. Workman et al. [32] argued that samples with high-resolution labels can be used to guide the training process in supervised learning using low-resolution labels. Region aggregation methods have also been used to improve the resolution of images. Quan et al. [33] utilized multi-scale edge features obtained by Differential Difference of Gaussian (DoG) methods to improve the results of edge extraction. Li et al. [34] proposed a progressive recurrent neural network to remove RS images destriping. Zhong et al. [35] proposed a transformer-based noise identification network model to help mitigate the over-segmentation phenomenon. However, semantic segmentation algorithms rely heavily on a large amount of training data, and it is difficult for these frameworks to achieve good results when labeled data is limited.
@@ II-C Semi-Supervised Semantic Segmentation in RS Images
The second category focuses on enhancements targeting domain-specific challenges in RS. For instance, Gao et al. [51] developed a semi-supervised road extraction network via multiple consistency and multitask constraints (MCMCNet). It ensures enhanced preservation of road continuity and trunk integrity. Chen et al. [52] proposed an adversarial learning-based semi-supervised network containing dual discriminators for road extraction from RS images. The two discriminators reinforce feature distribution alignment while maintaining consistency between labeled and unlabeled data. However, these semi-supervised methods typically focus on particular remote classes such as roads and traffic signs, demonstrating limited generalization. For the more widely used multi-class RS images, Xin et al. [8] proposed a hybrid framework integrating consistency regularization and contrastive learning to address interclass similarities and intraclass variations in RS imagery. Nevertheless, their experiments predominantly focused on a 25% labeled data ratio, failing to demonstrate effectiveness under very few labeled data scenarios. Huang et al.[53] considered the significant domain shift between different RS datasets. They proposed the semi-supervised bidirectional alignment for RS cross-domain scene classification. To address the inevitable issue of erroneous pseudo-labels and long-tail distribution in RS images, Huang et al.[9] introduced a Decoupled Weighting Learning (DWL) framework. During training, the decoupled learning module separates the predictions of labeled and unlabeled data. And the ranking weighting module attempts to adaptively weight each pseudo-label of the unlabeled data based on its relative confidence rank within the pseudo-class.
@@ III-A Main Optimization Objectives
We define $D^{L}=\{(x_{i}^{l},y_{i})\}_{i=1}^{N_{L}}$ as labeled data and $D^{U}=\{(x_{i}^{u})\}_{i=1}^{N_{U}}$ as unlabeled. Here ${x_{i}^{l}}\in\mathbb{R}^{H\times W\times 3}$ denotes the labeled image, ${y_{i}}\in\mathbb{R}^{H\times W\times K}$ is the ground truth of $K$ classes, while ${x_{i}^{u}}\in\mathbb{R}^{H\times W\times 3}$ denotes the unlabeled image, $N_{L}$ and $N_{U}$ are the amount of labeled and unlabeled images. $H$ and $W$ specify the height and width of the image. It is worth noting that in general $N_{U}$ is much larger than $N_{L}$. For unlabeled images, weak augmentation (WA) and strong augmentation (SA) are performed to train the student and teacher networks respectively. The main loss function is:
@@ IV-D Ablation study
MSUC encodes visual features from the four stages of the encoder, i.e., $V_{i}\in{\mathbb{R}^{C_{i}\texttimes H_{i}\texttimes W_{i}}}$, $i\in\{1,2,3,4\}$. Therefore, it is critical to explore the impact of multi-level (i.e., low to high) deep features on our MSUC. From Table III, we can see that the deep MSUC feature $V_{4}$ achieves greater improvement than the other low-level features, while the lowest feature $V_{1}$ makes the worst mIoU. The reason is that the deep features have a rich representation of information.
@@
Analysis of features fusion for MSUC module.
We fuse deep and low features stage by stage for the MSUC module. Specifically, we firstly employ the deep feature $V_{4}$ alone, and then gradually add all the other level features. Table IV shows the results of these different combinations. From the results, we observe that the mIoU gradually increases as lower level features are added stage by stage. This is because ̵low-level features provide small-scale structural details, while deep features capture large-scale contextual semantics. So, low level features contain more local information, which can effectively compensate deep features. The fusion of deep and low features improves the consistency of the model at different layers and scales, and optimizes the semantic segmentation results of RS images.
@@ IV-E Comparison with SOTA methods on RS Datasets
This section conducts experiments on ISPRS-Potsdam and LoveDA datasets, compared to the SOTA methods, including Mean teacher [10], CutMix [59], CCT [42], CPS [43], LSST [48], Fixmatch [11], Unimatch [18], DWL [9], and Allspark [12]. Specifically, we show the results for labeled data ratios of $1\%$, $5\%$ and $10\%$ with SegFormer-B2 as the segmentation model to verify the effectiveness, respectively. We also execute the experiments of OnlySup and FullySup for reference, where OnlySup only uses the labeled data for training and FullySup uses all data. The experimental results are presented in Tables V and VI, respectively.
@@ IV-G Visualization Comparison
We conducted visual comparison experiments to see the advantages of our approach more intuitively and clearly. Fig. 7 and Fig. 8 show the visual comparison of several semi-supervised semantic segmentation methods on the ISPRS-Potsdam dataset and LoveDA dataset.
For ISPRS-Potsdam dataset, the Fixmatch and Unimatch make errors in segmenting regions for the Low vegetation, Tree, Impervious surfaces and Building classes. Allspark and DWL successfully recognizes part of the region where Low vegetation classes are mixed with Tree classes, but the segmentation results are significantly enlarged for building classes. In contrast, our method shows significant advantages in accurately recognizing and segmenting Building (blue), Car (yellow), and Tree (green). Especially, we can see that methods including Fixmatch and Unimatch show large-scale segmentation errors when Car, Low vegetation and Tree are mixed. In addition, we overlap our segmentation results with the original image in the last column to clearly show our advantage. This overlapped image demonstrates that our model achieves excellent visual results, particularly in the Tree and Car categories.

### v3

@@ II-A RS semantic segmentation
Semantic segmentation is to segment an image into regional blocks with certain semantic meanings [22, 23], and recognize the specific semantic category of each regional block respectively. It commonly labels the image pixel by pixel [24, 25, 26, 27, 28]. Early semantic segmentation models including FCN [29], SegNet [30], and U-Net [31], achieved remarkable performance in traditional semantic segmentation tasks. With the success of Transformers in vision in recent years, more and more semantic segmentation models are adopting the Transformers architecture[32]. Complex backgrounds as well as resolution variations are the main challenges in semantic segmentation of RS images. In RS images, large intra-class and small inter-class variations between objects make image feature characterization difficult. Workman et al. [33] argued that samples with high-resolution labels can be used to guide the training process in supervised learning using low-resolution labels. Region aggregation methods have also been used to improve the resolution of images. Quan et al. [34] utilized multi-scale edge features obtained by Differential Difference of Gaussian (DoG) methods to improve the results of edge extraction. Li et al. [35] proposed a progressive recurrent neural network to remove RS images destriping. Zhong et al. [36] proposed a transformer-based noise identification network model to help mitigate the over-segmentation phenomenon. However, semantic segmentation algorithms rely heavily on a large amount of training data, and it is difficult for these frameworks to achieve good results when labeled data is limited.
@@ II-C Semi-Supervised Semantic Segmentation in RS Images
The second category focuses on enhancements targeting domain-specific challenges in RS. For instance, Gao et al. [52] developed a semi-supervised road extraction network via multiple consistency and multitask constraints (MCMCNet). It ensures enhanced preservation of road continuity and trunk integrity. Chen et al. [53] proposed an adversarial learning-based semi-supervised network containing dual discriminators for road extraction from RS images. The two discriminators reinforce feature distribution alignment while maintaining consistency between labeled and unlabeled data. However, these semi-supervised methods typically focus on particular remote classes such as roads and traffic signs, demonstrating limited generalization. For the more widely used multi-class RS images, Xin et al. [8] proposed a hybrid framework integrating consistency regularization and contrastive learning to address interclass similarities and intraclass variations in RS imagery. Nevertheless, their experiments predominantly focused on a 25% labeled data ratio, failing to demonstrate effectiveness under very few labeled data scenarios. Liu et al. [54] propose a Multi-Level Label-Aware (MLLA) semi-supervised framework for remote sensing scene classification to address inter-class similarities. By extending semantic information extraction from unlabeled data from single-level to multi-level, MLLA enhances scene classification performance. Zhang et al.[55] propose the MDMASNet framework to tackle the rich multi-scale information in remote sensing images. ̵MDMASNet employs a pyramid architecture that integrates deformable convolution and dilated convolution to extract features from objects of diverse shapes and sizes. Although these networks target rich multi-scale information and high inter-class similarities in semi-supervised remote sensing, their proposed network models ̵exhibit excessive complexity and fail to achieve non-intrusiveness. Huang et al.[56] considered the significant domain shift between different RS datasets. They proposed the semi-supervised bidirectional alignment for RS cross-domain scene classification. To address the inevitable issue of erroneous pseudo-labels and long-tail distribution in RS images, Huang et al.[9] introduced a Decoupled Weighting Learning (DWL) framework. During training, the decoupled learning module separates the predictions of labeled and unlabeled data. And the ranking weighting module attempts to adaptively weight each pseudo-label of the unlabeled data based on its relative confidence rank within the pseudo-class.
@@ III-A Main Optimization Objectives
We define $D^{L}=\{(x_{i}^{l},y_{i})\}_{i=1}^{N_{L}}$ as labeled data and $D^{U}=\{(x_{i}^{u})\}_{i=1}^{N_{U}}$ as unlabeled. Here ${x_{i}^{l}}\in\mathbb{R}^{H\times W\times 3}$ denotes the labeled image, ${y_{i}}\in\mathbb{R}^{H\times W\times K}$ is the ground truth of $K$ classes, while ${x_{i}^{u}}\in\mathbb{R}^{H\times W\times 3}$ denotes the unlabeled image, $N_{L}$ and $N_{U}$ are the amount of labeled and unlabeled images. $H$ and $W$ specify the height and width of the image. It is worth noting that in general $N_{U}$ is much larger than $N_{L}$. For unlabeled images, weak augmentation (WA) and strong augmentation (SA) are performed to train the teacher and student networks respectively. The main loss function is:
@@ IV-D Ablation study
MSUC encodes visual features from the four stages of the encoder, i.e., $V_{i}\in{\mathbb{R}^{C_{i}\texttimes H_{i}\texttimes W_{i}}}$, $i\in\{1,2,3,4\}$. Therefore, it is critical to explore the impact of multi-level (i.e., low to high) deep features on our MSUC. From Table III, we can see that the deep MSUC feature $V_{4}$ achieves greater improvement than the other low-level features, while the lowest feature $V_{1}$ makes the worst mIoU. The reason is that when only utilizing features extracted from a single stage of the model, deep features are more effective than low-level features on forming global semantic representations such as object categories and overall contours.
@@
Analysis of features fusion for MSUC module
We fuse deep and low features stage by stage for the MSUC module. Specifically, we firstly employ the deep feature $V_{4}$ alone, and then gradually add all the other level features. Table IV shows the results of these different combinations. From the results, we observe that the mIoU gradually increases as lower level features are added stage by stage. This is because ̵low-level features, generated by the shallow stages of the network, have small receptive fields and primarily capture ̵local details such as image edges and textures. Deep features accumulate ̵large-scale contextual information through repeated downsampling and convolutional operations, forming ̵global semantic representations like object categories and overall contours. Therefore, the ̵high resolution and ̵detail retention capability of low-level features make them inherently suited for ̵fine-grained analysis. While the ̵lower resolution and ̵broader receptive fields of deep features enable them to excel in ̵large-scale semantic comprehension. The fusion of deep and low features improves the consistency of the semi-supervised model at different stages and scales, significantly improving the model’s capability to capture multi-scale characteristics of target objects. ̵Consequently, it optimizes semantic segmentation performance in RS images.
@@ IV-E Comparison with SOTA methods on RS Datasets
This section conducts experiments on ISPRS-Potsdam and LoveDA datasets, compared to the SOTA methods, including Mean teacher [10], CutMix [62], CCT [43], CPS [44], LSST [49], Fixmatch [11], Unimatch [18], DWL [9], and Allspark [12]. Specifically, we show the results for labeled data ratios of $1\%$, $5\%$ and $10\%$ with SegFormer-B2 as the segmentation model to verify the effectiveness, respectively. We also conduct experiments of OnlySup and FullySup as reference, where OnlySup uses only 1%, 5%, and 10% of labeled data for training, while FullySup employs the entire dataset which represents the 100% labeled data. The experimental results are presented in Tables V and VI, respectively.
@@ IV-G Visualization Comparison
We conducted visual comparison experiments to see the advantages of our approach more intuitively and clearly. Fig. 7 and Fig. 8 show the visual comparison of several semi-supervised semantic segmentation methods on the ISPRS-Potsdam dataset and LoveDA dataset. Additionally, we have highlighted the mis-segmented areas of other SOTA methods with dashed ellipses, while emphasizing the comparatively accurate segmentation results of our approach in the same regions for contrast.
For ISPRS-Potsdam dataset, the Fixmatch and Unimatch make errors in segmenting regions for the Low vegetation, Tree, Impervious surfaces and Building classes. Allspark and DWL successfully recognize part of the region where Low vegetation classes are mixed with Tree classes, but the segmentation results for Building class exhibit inaccuracies. In contrast, our method shows significant advantages in accurately recognizing and segmenting Building (blue), Car (yellow), and Tree (green). Especially, we can see that methods including Fixmatch and Unimatch show large-scale segmentation errors when Car, Low vegetation and Tree are mixed. In addition, we overlap our segmentation results with the original image in the last column to clearly show our advantage. This overlapped image demonstrates that our model achieves excellent visual results, particularly in the Tree and Car categories.